OpenStudy (anonymous):

Trigonometry. Rationalize the numerator. Assume that all radicands are nonnegative. 3+sin y/3-sin y (all under one square root)

OpenStudy (anonymous):

thinking of multiplying both numerator and denominator by 3+siny? then change sin^2 to 1-cos^2, dunno

OpenStudy (anonymous):

ohh sqrt() did not see that, let me write it down!

OpenStudy (anonymous):

anyone helping?

OpenStudy (anonymous):

trying to find some paper grr

OpenStudy (phi):

Is this the problem \[ \sqrt{\frac{3+sin(y)}{3-sin(y))} } \]

OpenStudy (anonymous):

yes that's how it looks

OpenStudy (phi):

Rationalize the numerator (normally people rationalize the denominator) multiply top and bottom by sqrt(3+sin(y))

OpenStudy (anonymous):

doing that, but i get to 3+siny/sqrt(cos^2x+8)

OpenStudy (anonymous):

hmmm, think u need to do it like sqrt\[((\sqrt(3+sinx)/\sqrt(3-sinx))*(\sqrt(3-sinx)/\sqrt(3-sinx))\]

OpenStudy (anonymous):

so then is it sqrt() 9+sin^2/9-sin^2y ???

OpenStudy (anonymous):

so u get to \[\sqrt(9-\sin^2 x)/(3-\sin x)\]

OpenStudy (anonymous):

that is it! this is the answer, no square root at the denominator